### v1

# Achieving 70 Gb/s Over A VCSEL-Based Optical Wireless Link Using A Multi-Mode Fiber-Coupled Receiver

Abstract
In this paper, we demonstrate a laser-based optical wireless communication (OWC) system employing a nm single-mode (SM) vertical cavity surface emitting laser (VCSEL) and a multi-mode (MM) fiber-coupled receiver, achieving a record data rate beyond Gb/s, while the optical transmit power is below mW. The use of a high speed fiber-optic photoreceiver avoids limiting the communication bandwidth by the receiver, enabling ultra-high capacity and energy-efficient light fidelity (LiFi) links to unlock new applications. This work experimentally validates the feasibility of ultra-high speed indoor OWC systems using a single low-power and low-cost VCSEL for next-generation LiFi connectivity.
Index Terms:
Next-generation light fidelity (LiFi), laser-based optical wireless communication (OWC), ultra-high speed, vertical cavity surface emitting laser (VCSEL), fiber-optic photoreceiver.I Introduction
The next generation of light fidelity (LiFi) technology is set to leverage laser-based optical wireless communication (OWC) to overcome the bandwidth limitations of conventional light-emitting diode (LED)-based OWC systems and meet the stringent requirements of sixth generation (6G) networks [1]. Laser-based OWC offers a cost-effective solution by enabling ultra-high speed and energy-efficient wireless links to support real-time bandwidth-intensive applications in 6G targeting multi-Gb/s data rates with ultra-low latency, such as augmented reality (AR), virtual reality (VR), 4K and 8K ultra-high-definition (UHD) video streaming, ultra-high data density services, and inter-rack connections in data centers [1, 2, 3].
Achieving data rates of tens of Gb/s using laser-based OWC for indoor applications has been actively investigated [4, 5, 6]. In [4], Lee et al. experimentally validated a dual‑wavelength wavelength division multiplexing (WDM) OWC system using a novel white‑light transmitter by integrating a nm gallium nitride (GaN) blue laser, an nm infrared (IR) laser, and a phosphor converter within a single surface‑mount device (SMD) package. The authors reported a combined data rate exceeding Gb/s over a m link at a target bit error ratio (BER) of adopted from long term evolution (LTE) systems, using two Silicon positive-intrinsic-negative (PIN) photoreceivers with a dB bandwidth of GHz to detect the optical signals of the blue and IR laser diodes. In [5], Chen et al. implemented a ‑channel LiFi link using the same SMD laser technology based on a transmitter setup that integrates ten high‑brightness, high‑bandwidth LDs, including three GaN blue LDs operating at nm, nm, and nm, and seven IR LDs with wavelegths of nm to nm. To detect the optical signals, GHz Silicon PIN photoreceivers were used, except for the nm LD, for which a GHz indium gallium arsenide (InGaAs) photoreceiver was used due to its higher responsivity. The authors demonstrated aggregate data rates in excess of Gb/s over a few meter link distance by considering a BER threshold of .
Among various types of LDs, vertical cavity surface emitting lasers are well-suited for realizing laser-based OWC due to their outstanding features. They are low-cost and low-power devices (e.g., with a few mW power consumption), and have an extensive modulation bandwidth (e.g., GHz), well-controlled output beam properties, and the ability to be fabricated as two-dimensional (2D) arrays [7]. In particular, single-mode (SM) VCSELs emit a Gaussian laser beam with a narrow spectral linewidth [8]. In addition, SM VCSELs offer a lower far field (FF) divergence angle compared to their multi-mode (MM) counterparts, providing a smaller beam spot size which leads to improved light coupling efficiency when the beam is focused on the small active area of a high speed photodiode (PD). This is of course an advantage for high speed point-to-point optical wireless link design. In [6], Osahon et al. experimentally demonstrated a line‑of‑sight optical wireless link based on an nm SM VCSEL with a -dB bandwidth of GHz, achieving a gross data rate of Gb/s over m with an eye‑safe launch power of dBm (i.e., mW), assuming a forward error correction (FEC) limit for a pre-FEC BER of . This performance was achieved by using a GHz Silicon PIN PD and spreading the data across a large number of discrete multitone (DMT) subcarriers by means of adaptive bit and power loading.
In this work, we demonstrate a laser-based optical wireless link experiment using a nm SM VCSEL, reaching data rates of Gb/s with an average optical power of mW. The VCSEL has an ultra-wide modulation bandwidth in excess of GHz. In order to realize the full potential of this VCSEL, we employ a GHz fiber-coupled receiver to ensure that the receiver does not limit the end-to-end bandwidth of the system. In addition, we apply direct current-biased optical orthogonal frequency division multiplexing (DCO-OFDM) with adaptive bit and power loading. To the best of our knowledge, this is the first experimental proof of a directly modulated SM VCSEL at nm in conjunction with a MM fiber-optic photoreceiver for ultra-high speed laser-based OWC, achieving a record data rate of Gb/s using a single low-power VCSEL. This brings about a fourfold increase in the achievable data rate compared to our previous work [9].
The rest of the paper is organized as follows. In Section II, the experimental setup of the proposed VCSEL-based OWC system is described. In Section III, the measurement results are discussed, in terms of the frequency response of the end-to-end system, the adaptive bit and power loading performance, the achievable data rate and the BER performance. In Section IV, concluding remarks and future works are presented.
II Experimental Setup
II-A End-to-End System Setup
Fig. 1 shows the link configuration including the optical and optomechanical assemblies of the transmitter and receiver. We use an SM VCSEL operating at a center wavelength of nm with a narrow linewidth. This VCSEL was designed, manufactured and packaged by Vertically Integrated Systems (VIS) GmbH and was tested at the wafer level before being fully characterized for optical communications [10]. It is a multi-aperture VCSEL with a compact mini-array structure, comprising four VCSELs that are arranged in a array configuration with a small pitch of nm111Such a miniature structure is fabricated using oxidation through holes [10]. [10]. These four VCSELs are connected to a common pair of contacts such that they act as a single VCSEL. As a result, the output power is effectively increased while maintaining the SM emission spectrum [10]. The VCSEL mini-array has a nearly Gaussian beam profile with a FF full-width at half-maximum (FWHM) divergence angle of , a maximum optical power of mW, and a -dB modulation bandwidth of GHz. Furthermore, it provides a linear input dynamic range between its forward threshold current of mA and roll-over current of mA, which is suitable for DCO-OFDM modulation as discussed in the following.
Fig. 2 depicts the complete experimental setup. In this work, considering the low-pass and frequency-selective characteristics of the VCSEL-based OWC system, we use DCO-OFDM with adaptive subcarrier bit and power allocation to ensure high spectral efficiency. The modulation and demodulation of the OFDM signal are implemented based on -point inverse fast Fourier transform (IFFT) and FFT operations, respectively, where , using offline digital signal processing (DSP) in MATLAB running on a laptop computer. At the transmitter, the information bits are mapped to -ary quadrature amplitude modulation (QAM) symbols to be loaded across a total of data-carrying subcarriers, with the constellation size reaching up to , depending on the available signal-to-noise ratio (SNR). The DCO-OFDM frame is constructed by concatenating the block of the complex QAM symbols with its conjugated version according to a Hermitian symmetry. A cyclic prefix (CP) of length is then added and an IFFT is performed on the resulting frame to generate the time domain OFDM signal. This signal is further shaped using a root-raised cosine filter with a roll-off factor of 0.1. The pulse-shaped digital waveform is converted into a wideband analog signal via the Keysight M8195A arbitrary waveform generator (AWG), operating at a sampling frequency of GSa/s. According to the Nyquist-Shannon theorem, the bandwidth of the signal is related to the sampling frequency via . In practice, the signal bandwidth at the AWG output can be controlled using the number of samples per symbol (SPS) parameter, denoted by , based on . We choose to obtain GHz. The AWG output signal is combined with the required direct current (DC) bias using a Mini Circuits ZX85-12G-S+ bias tee (BT) to drive the VCSEL. This BT has a GHz bandwidth with a minimum operating frequency of kHz, and an insertion loss (IL) of dB. The wideband OFDM signal is fed into the BT using a radio frequency (RF) coaxial cable with a bandwidth of GHz, a length of cm, and dB over the desired bandwidth. To collimate the light beam at the transmitter, a Thorlabs ACL5040U-B aspheric condenser lens is placed in front of the VCSEL. This lens has a diameter of mm with a clear aperture (CA) of mm, a focal length of mm, a numerical aperture (NA) of , and an anti-reflective coating (ARC) for – nm.
At the receiver side, a Thorlabs C80FC-B achromatic fiber collimator is employed to collect the incident light beam from the transmitter located at a cm distance. This collimator has a CA of mm, an effective focal length of mm, an ARC for – nm matching the coating of the transmitter lens, and a Ferrule Connector/Physical Contact (FC/PC) fiber-optic output with a maximum fiber NA of . As shown in Fig. 1, the fiber collimator is mounted on a three-axis kinematic mount with high-resolution angular adjusters along with an coordinates translation stage, providing six degrees of freedom required for precise alignment of the focused beam spot on the fiber port to ensure high coupling efficiency. The collimator output is connected to the optical input of the MM fiber-coupled receiver using an OM4 graded-index fiber patch cable of length m with a core diameter of µm and FC/PC connectors. We use a Thorlabs RXM25BF fiber-optic photoreceiver, which is equipped with an InGaAs PIN PD followed by integrated transimpedance amplifier (TIA) and variable gain amplifier (VGA) stages. This receiver has a broadband responsivity in – nm, and a photodetection bandwidth of up to GHz with a minimum operating frequency of kHz. The built-in VGA helps with adjusting the transimpedance gain and the peak of the receiver frequency response to optimize the performance. There is also a digital screen on the photoreceiver to monitor the average detected photocurrent, which helps with performing the alignment using an unmodulated laser beam. The output analog signal from the photoreceiver is converted to a digital signal by the Keysight UXR0104B oscilloscope at a sampling frequency of GSa/s. The remaining DSP is performed in MATLAB. This includes frame synchronization, FFT, single-tap equalization, and QAM demapping. For each experiment, a total of OFDM frames consisting of pilot frames and data frames are sent through the OWC system to measure the BER performance.
II-B Achievable Data Rate
In the proposed setup, the system components are carefully selected to support wideband signal transmission. In order to evaluate the system bandwidth, we need to take into account every single component or connection along the RF signal path in the end-to-end system. The system bandwidth is limited by the one that has the lowest bandwidth. However, the use of adaptive bit and power loading allows for the signal bandwidth to be extended beyond the -dB bandwidth of the system, due to the allocation of bits to high-frequency subcarriers with low SNR values. The data rate of the DCO-OFDM system is computed by [11]:
where is the number of bits allocated to the th subcarrier such that for or depending on the available SNR, with denoting the QAM constellation size for the th subcarrier. To meet a given BER requirement, the maximum value of is obtained as [12]:
where represents the corresponding subcarrier SNR, and is defined as the SNR gap, which is given by [12]:
Config. I | V | mA | µW |
Config. II | V | mA | µW |
III Measurement Results
We present the measurement results for the performance of the VCSEL-based OWC system in terms of the received SNR, average BER and link data rate. The channel transfer function is estimated by sending pilot -QAM symbols of unit variance on all data-carrying subcarriers [5]. The estimated channel information is then used to determine the optimal average transmit power and constellation order of QAM symbols for each subcarrier based on the corresponding subcarrier SNR, thus adapting the transmission rate and maximizing the system spectral efficiency. In this work, the Hughes-Hartogs (HH) bit and power loading algorithm is applied [13], which is widely used in multi-carrier communication systems. The optimal DC operating point of the VCSEL is obtained through extensive measurements and two configurations are identified. These configurations along with the resulting average received optical powers, denoted by , are given in Table I. In addition, to ensure operation within the linear dynamic range of the VCSEL, the peak-to-peak amplitude (i.e., standard deviation) of the OFDM signal is set to V. This value is acquired by sweeping in an appropriate range according to the light-current-voltage (LIV) characteristics of the VCSEL and specifying the value that maximizes the link data rate by post-processing the measurement data in MATLAB.
[…]
The estimated channel gain and SNR, as shown in Fig. 3, are used to allocate the optimal number of bits and amount of power to each subcarrier. Fig. 4 shows an example of adaptive bit loading and power allocation results. It can be observed that the number of loaded bits is greatly dependent on the estimated SNR across the operating bandwidth. Also, adaptive power allocation balances out the SNR level to ensure that it is adequate to send the allocated number of bits at the required BER. It can be observed that to bits are loaded on low frequency subcarriers up to GHz, to bits are loaded to high frequency subcarriers between GHz and GHz, and to bits are carried by the remaining subcarriers between GHz and GHz.
Fig. 5 shows the average BER performance with respect to the achievable data rate. It can be observed that the achievable data rate increases steadily at the cost of increasing the BER at the same time. It can also be observed that Config. I performs slightly better than Config. II. A possible explanation for this is Config. I sets the VCSEL’s DC bias at a lower operating point with slightly smaller voltage and current values. Therefore, the modulating signal experiences less nonlinear distortion effects, particularly for the signal swing above the DC operating point. In other words, in Config. I, the OFDM signal variations match the linear dynamic range of the VCSEL. As a result, the SNR is slightly improved especially in the lower frequency region up to GHz, as shown in Fig. 3, thereby improving the bit loading performance, as shown in Fig. 4. We consider a target BER of according to system-level requirements in LTE systems. In fact, in LTE systems, link adaptation aims to maintain a block error rate (BLER) of after decoding [14, Chapter 10, pp. 215–248]. This allows the system to tolerate a relatively high raw, uncoded BER on the order of to at the physical layer depending on the selected modulation and coding scheme (MCS) and the operating SNR so as to maximize spectral efficiency, relying on advanced FEC coding schemes such as Turbo codes or hybrid automatic repeat request (HARQ) mechanisms to enhance the BER performance [15]. The final post-FEC BER is expected to be at or lower after decoding with to overhead [16]. The BER performance can be further reduced to after decoding by using staircase codes with a overhead [16]. From Fig. 5, it can be observed that the achievable data rate exceeds Gb/s while . More precisely, the VCSEL-based optical wireless link achieves data rates of Gb/s and Gb/s at and under Config. I and Config. II, respectively. We note that both configurations achieve a data rate above Gb/s at a pre-FEC BER level of , which offers a improvement relative to the data rate achieved in [6] at the same BER level.
IV Conclusions
In this paper, we have demonstrated an experimental proof-of-concept for ultra-high data rate laser-based optical wireless links for future bandwidth-intensive applications. To this end, we employ a multi-aperture SM VCSEL operating at nm with an ultra-wide modulation bandwidth of GHz. We also employ a GHz MM fiber-optic photoreceiver to allow the VCSEL’s modulation bandwidth to be utilized by applying DCO-OFDM with adaptive subcarrier bit and power loading. The experimental results corroborate that the proposed system setup achieves a record-high data rate of Gb/s at a pre-FEC BER threshold of which is below the LTE target (i.e., ), while it attains Gb/s at a pre-FEC BER threshold of , over a link distance of cm. The link range and the received SNR can be increased using meticulous alignment procedures to maximize optical coupling efficiency at the MM fiber port of the receiver collimator. Furthermore, the GHz cutoff frequency of the oscilloscope is currently limiting the end-to-end performance of the system. Otherwise, considering the VCSEL bandwidth and the SNR response, this system can achieve beyod Gb/s. Therefore, more work is yet to be done to realize the full potential of the SM VCSEL for ultra-high speed OWC system design.
Acknowledgement
This work was financially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grant EP/Y037243/1 ‘Platform Driving The Ultimate Connectivity (TITAN)’.