### v2

# Achieving 70 Gb/s Over A VCSEL-Based Optical Wireless Link Using A Multi-Mode Fiber-Coupled Receiver

Abstract
In this paper, we demonstrate a laser-based optical wireless communication (OWC) system employing a nm single-mode (SM) vertical cavity surface emitting laser (VCSEL) and a multi-mode (MM) fiber-coupled receiver, achieving a record data rate beyond Gb/s, while the optical transmit power is below mW. The use of a high speed fiber-optic photoreceiver avoids limiting the communication bandwidth by the receiver, enabling ultra-high capacity and energy-efficient light fidelity (LiFi) links to unlock new applications. This work experimentally validates the feasibility of ultra-high speed indoor OWC systems using a single low-power and low-cost VCSEL for next-generation LiFi connectivity.
I Introduction
The next generation of light fidelity (LiFi) technology is set to leverage laser-based optical wireless communication (OWC) to overcome the bandwidth limitations of conventional light-emitting diode (LED)-based OWC systems and meet the stringent requirements of sixth generation (6G) networks [1]. Laser-based OWC offers a cost-effective solution by enabling ultra-high speed and energy-efficient wireless links to support real-time bandwidth-intensive applications in 6G targeting multi-Gb/s data rates with ultra-low latency, such as augmented reality (AR), virtual reality (VR), 4K and 8K ultra-high-definition (UHD) video streaming, ultra-high data density services, and inter-rack connections in data centers [1, 2, 3].
Achieving data rates of tens of Gb/s using laser-based OWC for indoor applications has been actively investigated [4, 5, 6]. In [4], Lee et al. experimentally validated a dual‑wavelength wavelength division multiplexing (WDM) OWC system using a novel white‑light transmitter by integrating a nm gallium nitride (GaN) blue laser, an nm infrared (IR) laser, and a phosphor converter within a single surface‑mount device (SMD) package. The authors reported a combined data rate exceeding Gb/s over a m link at a target bit error ratio (BER) of adopted from long term evolution (LTE) systems, using two silicon (Si) positive-intrinsic-negative (PIN) photoreceivers with a dB bandwidth of GHz to detect the optical signals of the blue and IR laser diodes. In [5], Chen et al. implemented a ‑channel LiFi link using the same SMD laser technology based on a transmitter setup that integrates ten high‑brightness, high‑bandwidth LDs, including three GaN blue LDs operating at nm, nm, and nm, and seven IR LDs with wavelegths of nm to nm. To detect the optical signals, GHz Si PIN photoreceivers were used, except for the nm LD, for which a GHz indium gallium arsenide (InGaAs) photoreceiver was used due to its higher responsivity. The authors demonstrated aggregate data rates in excess of Gb/s over a few meter link distance by considering a BER threshold of .
[…]
In this work, we demonstrate a laser-based optical wireless link experiment using a nm SM VCSEL, reaching data rates of Gb/s with an average optical power of mW. The VCSEL has an ultra-wide modulation bandwidth in excess of GHz. In order to realize the full potential of this VCSEL, we employ a GHz fiber-coupled receiver to ensure that the receiver does not limit the end-to-end bandwidth of the system. In addition, we apply direct current-biased optical orthogonal frequency division multiplexing (DCO-OFDM) with adaptive bit and power loading. To the best of our knowledge, this is the first experimental proof of a directly modulated SM VCSEL at nm in conjunction with a MM fiber-optic photoreceiver for ultra-high speed laser-based OWC, achieving a record data rate of Gb/s using a single low-power VCSEL. This brings about a fourfold increase in the achievable data rate compared to our previous work [12].
The rest of the paper is organized as follows. In Section II, the experimental setup of the proposed VCSEL-based OWC system is described. In Section III, the measurement results are discussed, in terms of the frequency response of the end-to-end system, the adaptive bit and power loading performance, the achievable data rate and the BER performance. In Section IV, concluding remarks and future works are presented.
II Experimental Setup
II-A VCSEL Characterization
We use an SM VCSEL operating at a center wavelength of nm with a low spectral width. This VCSEL was designed, manufactured and packaged by Vertically Integrated Systems (VIS) GmbH and was tested at the wafer level before being fully characterized for optical communications [13]. It is a multi-aperture (MA) VCSEL featuring a compact mini-array structure, comprising four oxide-confined VCSELs, each with an effective aperture diameter of about µm, which are arranged in a array configuration with a small pitch of µm [13]. These four VCSELs have a common pair of contacts such that they act as a single VCSEL. Consequently, the output power of the MA VCSEL is effectively improved by a factor of four while maintaining both the SM emission spectrum and the broadband modulation response of individual VCSELs in the mini-array structure [13]. The small-pitch VCSEL mini-array has a compact footprint for efficient light coupling into a µm core MMF [7]. Such a miniature structure is fabricated using oxidation through holes [7]. A die shot of the VCSEL mini-array chip is available in [13].
Fig. 1 shows the light-current-voltage (LIV) characteristics of the VCSEL measured at room temperature. The VCSEL has a maximum output power of mW, a forward threshold current of about mA, and a roll-over current of mA. In addition, it provides a linear input dynamic range with a slope efficiency of W/A for forward currents between mA and mA, which is suitable for DCO-OFDM modulation. Fig. 2 shows the optical spectrum of the VCSEL for different values of the direct current (DC) bias current , measured using a high-resolution optical spectrum analyzer (OSA). It can be observed that when increases from mA to mA, the peak wavelength shifts slightly from nm to nm. This spectral shift is mainly due to thermal effects, which cause the longitudinal resonance of the cavity to move to longer wavelengths [7]. For each case, the average wavelength and the root mean square (RMS) spectral width are also provided in Fig. 2. The RMS spectral width is a measure of the effective wavelength spread of the optical spectrum and is defined as the weighted standard deviation of the wavelength around its mean [14]:
where is the emitted power at wavelength . The value of directly determines the chromatic dispersion (CD) penalty and maximum reachable distance in MMF links [7]. According to Fig. 2, by increasing from mA to mA, increases from nm to nm. As a reference for comparison, the IEEE 802.3 standard for Ethernet over MMF links specifies a maximum RMS spectral width of nm for VCSELs operating at nm to nm to reach link distances up to m [15]. Fig. 3 shows the FF intensity profile of the VCSEL measured at different values of , which confirms that the VCSEL has a nearly Gaussian beam profile, with a full-width at half-maximum (FWHM) divergence angle ranging between and when increasing from mA to mA. The corresponding divergence half-angle, defined for the beam spot containing of the transmitted power, varies between and . Furthermore, the VCSEL has a -dB modulation bandwidth of GHz.
II-B End-to-End System Setup
Fig. 4 shows the link configuration including the optical and optomechanical assemblies of the transmitter and receiver. Fig. 5 depicts the complete experimental setup. In this work, considering the low-pass and frequency-selective characteristics of the VCSEL-based OWC system, we use DCO-OFDM with adaptive subcarrier bit and power allocation to ensure high spectral efficiency. The modulation and demodulation of the OFDM signal are implemented based on -point inverse fast Fourier transform (IFFT) and FFT operations, respectively, where , using offline digital signal processing (DSP) in MATLAB running on a laptop computer. At the transmitter, the information bits are mapped to -ary quadrature amplitude modulation (QAM) symbols to be loaded across a total of data-carrying subcarriers, with the constellation size reaching up to , depending on the available signal-to-noise ratio (SNR). The DCO-OFDM frame is constructed by concatenating the block of the complex QAM symbols with its conjugated version according to a Hermitian symmetry. A cyclic prefix (CP) of length is then added and an IFFT is performed on the resulting frame to generate the time domain OFDM signal. This signal is further shaped using a root-raised cosine filter with a roll-off factor of 0.1. The pulse-shaped digital waveform is converted into a wideband analog signal via the Keysight M8195A arbitrary waveform generator (AWG), operating at a sampling frequency of GSa/s. According to the Nyquist-Shannon theorem, the bandwidth of the signal is related to the sampling frequency via . In practice, the signal bandwidth at the AWG output can be controlled using the number of samples per symbol (SPS) parameter, denoted by , based on . We choose to obtain GHz. The AWG output signal is combined with the required DC bias using a Mini Circuits ZX85-12G-S+ bias tee (BT) to drive the VCSEL. This BT has a GHz bandwidth with a minimum operating frequency of kHz, and an insertion loss (IL) of dB. The wideband OFDM signal is fed into the BT using a radio frequency (RF) coaxial cable with a bandwidth of GHz, a length of cm, and dB over the desired bandwidth. To collimate the light beam at the transmitter, a Thorlabs ACL5040U-B aspheric condenser lens is placed in front of the VCSEL. This lens has a diameter of mm with a clear aperture (CA) of mm, a focal length of mm, a numerical aperture (NA) of , and an anti-reflective coating (ARC) for – nm. Note that when the signal is directly modulated on the forward current of the MA VCSEL for intensity modulation (IM), although the output beams of the four apertures are overlapping while propagating in free space, there is no crosstalk between them, since they carry the same signal and their signal contributions are constructively added together upon direct detection (DD) at the receiver.
At the receiver side, a Thorlabs C80FC-B achromatic fiber collimator (AFC) is employed to collect the incident beam from the transmitter located at a cm distance. This AFC has a CA of mm, an effective focal length of mm, an ARC for – nm matching the coating of the transmitter lens, and a Ferrule Connector/Physical Contact (FC/PC) fiber-optic output with a maximum fiber NA of . As shown in Fig. 4, the AFC is mounted on a three-axis kinematic mount with high-resolution angular adjusters along with an coordinates translation stage, providing six degrees of freedom (DOF) required for precise alignment of the focused beam spot on the fiber plane to ensure high coupling efficiency. The output port of the AFC is connected to the optical input of the fiber-coupled receiver using a graded-index OM4 MMF patch cable of length m with a core diameter of µm and FC/PC connectors. We use a Thorlabs RXM25BF fiber-optic photoreceiver, which is equipped with an InGaAs PIN PD followed by integrated transimpedance amplifier (TIA) and variable gain amplifier (VGA) stages. This receiver has a broadband responsivity in – nm, and a photodetection bandwidth of up to GHz with a minimum operating frequency of kHz. The built-in VGA helps with adjusting the transimpedance gain and the peak of the receiver frequency response to optimize the performance. There is also a digital screen on the photoreceiver to monitor the average detected photocurrent, which helps with performing the alignment using an unmodulated laser beam. The output analog signal from the photoreceiver is converted to a digital signal by the Keysight UXR0104B oscilloscope at a sampling frequency of GSa/s. The remaining DSP is performed in MATLAB. This includes frame synchronization, FFT, single-tap equalization, and QAM demapping. For each experiment, a total of OFDM frames consisting of pilot frames and data frames are sent through the OWC system to measure the BER performance. As shown in Fig. 4, the back-to-back reference signal provides a copy of the transmitted OFDM signal on the oscilloscope Channel 2 through a cable connection with the AWG, which can be used as the triggering channel to capture the waveform of the received OFDM signal.
Since the AFC is mounted by using an optomechanical assembly with six DOF, any misalignment including lateral, angular, or axial, has an impact on the light coupling into the MMF core. It is particularly sensitive to angular misalignments when adjusting the three tilt angles of the AFC. Even a tiny turn of one of the adjusting knobs on the kinematic mount can greatly shift the beam spot on the focal plane of the AFC, causing the coupling efficiency to drop dramatically, which makes the alignment process practically challenging. We choose a link distance of cm to alleviate the complexity of alignment and achieve reliable performance. However, the proposed setup can be systematically reconfigured to increase the link distance. Note that the end-to-end performance is not limited by fiber dispersion or attenuation due to the short length of the OM4 MMF cable. To elaborate, Appendix A provides the maximum bandwidth and the maximum reachable distance of MMF links under both CD and modal dispersion (MD). The fiber patch cable has an effective modal bandwidth (EMB) of MHzkm and a loss of dB/km at nm [16]. We adopt these values to present a conservative estimate of the performance at nm, since EMB increases and decreases with wavelength [17]. Based on (14), by substituting m, ps/nmkm [18], and nm from Fig. 2, we obtain THz. Compared to the signal bandwidth GHz, , indicating that CD and MD have practically no impact on the performance. Alternatively, based on (15), we find m, which leads to the same conclusion. In addition, the fiber attenuation for m is dB, which is quite negligible.
II-C Achievable Data Rate
In the proposed setup, the system components are carefully selected to support wideband signal transmission. In order to evaluate the system bandwidth, we need to take into account every single component or connection along the RF signal path in the end-to-end system. The system bandwidth is limited by the one that has the lowest bandwidth. However, the use of adaptive bit and power loading allows for the signal bandwidth to be extended beyond the -dB bandwidth of the system, due to the allocation of bits to high-frequency subcarriers with low SNR values. The data rate of the DCO-OFDM system is computed by [19]:
where is the number of bits allocated to the th subcarrier such that for or depending on the available SNR, with denoting the QAM constellation size for the th subcarrier. To meet a given BER requirement, the maximum value of is obtained as [20]:
where represents the corresponding subcarrier SNR, and is defined as the SNR gap, which is given by [20]:
Hence, the achievable data rate in (2) is upper-bounded as:
| [V] | [mA] | [mW] | [µW] | |
| Config. I | ||||
| Config. II |
: Transmitted optical power.
III Measurement Results
We present the measurement results for the performance of the VCSEL-based OWC system in terms of the received SNR, average BER and link data rate. The channel transfer function is estimated by sending pilot -QAM symbols of unit variance on all data-carrying subcarriers [5]. The estimated channel information is then used to determine the optimal average transmit power and constellation order of QAM symbols for each subcarrier based on the corresponding subcarrier SNR, thus adapting the transmission rate and maximizing the system spectral efficiency. In this work, the Hughes-Hartogs (HH) bit and power loading algorithm is applied [21], which is widely used in multi-carrier communication systems. The optimal DC operating point of the VCSEL is obtained through extensive measurements and two configurations are identified. These configurations along with the resulting average received optical powers, denoted by , are given in Table I. In addition, to ensure operation within the linear dynamic range of the VCSEL, the peak-to-peak amplitude (i.e., standard deviation) of the OFDM signal is set to V. This value is acquired by sweeping in an appropriate range according to the LIV characteristics of the VCSEL and specifying the value that maximizes the link data rate by post-processing the measurement data in MATLAB.
III-A SNR Frequency Response: Adaptive Bit and Power Loading
Fig. 6 shows the channel gain and the electrical SNR as a function of frequency. This channel gain represents the channel transfer function which is measured in the frequency domain between the IFFT input at the transmitter and the FFT output at the receiver. It can be observed that the channel gain exhibits a sharp decline after GHz. This is primarily caused by the hard cutoff frequency of the oscilloscope at about GHz. Although the VCSEL and the fiber-coupled receiver can reach a higher bandwidth (i.e., GHz and GHz, respectively), in this experiment, the oscilloscope limits the bandwidth of the end-to-end system. We note that since the BT has a bandwidth of GHz, it also causes a roll-off on the frequency response at high frequencies. However, the cutoff frequency at GHz is still dominant. Evidently, the resulting SNR values are above dB over the entire range of frequencies up to GHz. Thus, the effective modulation bandwidth that can be used for data transmission on OFDM subcarriers is GHz. It can be observed that for GHz, dB for GHz, and dB for GHz. Furthermore, there are oscillations in the SNR response especially at frequencies lower than GHz, due to the impedance mismatch between the VCSEL and the BT. In fact, while the BT has an output impedance of Ω, the input impedance of the VCSEL is significantly less than Ω (i.e., about Ω). Although the VCSEL chip is packaged on an standard RF-V connector and it is directly connected to the BT output, as shown in Fig. 4, its low input impedance leads to signal reflections at this critical connection, resulting in RF interference that degrades the SNR and hence oscillations appear in the SNR frequency response. The remaining cable connections along the RF signal path at the transmitter and receiver are carefully chosen to be as short as possible to minimize RF signal distortions.
The estimated channel gain and SNR, as shown in Fig. 6, are used to allocate the optimal number of bits and amount of power to each subcarrier. Fig. 7 shows an example of adaptive bit loading and power allocation results. It can be observed that the number of loaded bits is greatly dependent on the estimated SNR across the operating bandwidth. Also, adaptive power allocation balances out the SNR level to ensure that it is adequate to send the allocated number of bits at the required BER. It can be observed that to bits are loaded on low frequency subcarriers up to GHz, to bits are loaded to high frequency subcarriers between GHz and GHz, and to bits are carried by the remaining subcarriers between GHz and GHz.
III-B BER vs Achievable Data Rate
Fig. 8 shows the average BER performance with respect to the achievable data rate. It can be observed that the achievable data rate increases steadily at the cost of increasing the BER at the same time. It can also be observed that Config. I performs slightly better than Config. II. A possible explanation for this is Config. I sets the VCSEL’s DC bias at a lower operating point with slightly smaller voltage and current values. Therefore, the modulating signal experiences less nonlinear distortion effects, particularly for the signal swing above the DC operating point. In other words, in Config. I, the OFDM signal variations match the linear dynamic range of the VCSEL. As a result, the SNR is slightly improved especially in the lower frequency region up to GHz, as shown in Fig. 6, thereby improving the bit loading performance, as shown in Fig. 7. We consider a target BER of according to system-level requirements in LTE systems. In fact, in LTE systems, link adaptation aims to maintain a block error rate (BLER) of after decoding [22, Chapter 10, pp. 215–248]. This allows the system to tolerate a relatively high raw, uncoded BER on the order of to at the physical layer depending on the selected modulation and coding scheme (MCS) and the operating SNR so as to maximize spectral efficiency, relying on advanced FEC coding schemes such as Turbo codes or hybrid automatic repeat request (HARQ) mechanisms to enhance the BER performance [23]. The final post-FEC BER is expected to be at or lower after decoding with to overhead [24]. The BER performance can be further reduced to after decoding by using staircase codes with a overhead [24]. From Fig. 8, it can be observed that the achievable data rate exceeds Gb/s while . More precisely, the VCSEL-based optical wireless link achieves data rates of Gb/s and Gb/s at and under Config. I and Config. II, respectively. We note that both configurations achieve a data rate above Gb/s at a pre-FEC BER level of , which offers a improvement relative to the data rate reported in [6] at the same BER level.
III-C Achieving Gb/s Through SNR Extrapolation
Looking at Fig. 6, a natural question arises as to how much performance enhancement can be achieved if the oscilloscope bandwidth constraint is eliminated? To address this question, we estimate the trend of the SNR in the frequency domain by approximating the measured SNR frequency response based on piecewise linear (PWL) regression. For each configuration, the approximated SNR is characterized by two frequencies of and such that , where GHz indicates the frequency at which the SNR drops to dB, as shown in Fig. 9. A moving average filter with a window size of is also employed to smooth out the ripples in the measured SNR results due to RF reflections. The resulting curve manifests the SNR trend and is used as a guide for line fitting in each frequency band. The computed values of and for both configurations are listed in Table II. The approximated SNR is then extrapolated by removing the hard limit of the oscilloscope at and letting the SNR to follow the same decreasing trend as in the frequency band until it crosses dB at , as shown in Fig. 9. This leads to GHz for Config. I, and GHz for Config. II. The rationale behind this extrapolation lies in the fact that the frequency response of SM MA VCSELs exhibits a monotonically decreasing behavior with a consistent roll-off at high frequencies [25]. Thus, when there is no other factor limiting the end-to-end modulation bandwidth, the system frequency response follows a similar decreasing trend.
| [GHz] | [GHz] | [GHz] | |
| Config. I | |||
| Config. II |
For an additive white Gaussian noise (AWGN) channel with an SNR frequency response of , the achievable data rate can be upper-bounded as [26]:
where is given by (4). When evaluating the upper bound in (6) for the approximated and extrapolated SNR, we have:
where is replaced by for the approximated SNR, and by for the extrapolated SNR. Fig. 10 shows the BER performance with respect to the achievable data rate for both cases based on the upper bound in (7). As observed from Fig. 10, under the approximated SNR, the upper bound of the data rate gets tighter for higher BER values, indicating that it can be safely used to estimate the maximum achievable data rate. With the extrapolated SNR, the achievable data rate is remarkably enhanced, reaching a maximum value of Gb/s and Gb/s at a BER of and for Config. I and Config. II, respectively. This confirms that the system can achieve data rates Gb/s.
IV Conclusions
In this paper, we have demonstrated an experimental proof-of-concept for ultra-high data rate laser-based optical wireless links for future bandwidth-intensive applications. To this end, we employ a SM MA VCSEL operating at nm with an ultra-wide modulation bandwidth of GHz. We also employ a GHz MM fiber-optic photoreceiver to allow the VCSEL’s modulation bandwidth to be utilized by applying DCO-OFDM with adaptive subcarrier bit and power loading. The experimental results corroborate that the proposed system setup achieves a record-high data rate of Gb/s at a pre-FEC BER of which is below the LTE target (i.e., ), while it attains Gb/s at a pre-FEC BER of , over a link distance of cm. The link range and the received SNR can be increased using meticulous alignment procedures to maximize optical coupling efficiency at the MMF port of the receiver collimator. Furthermore, the GHz cutoff frequency of the oscilloscope is currently limiting the end-to-end performance of the system. Otherwise, considering the VCSEL bandwidth and the SNR response, this system can reach data rates beyond Gb/s. Through extrapolating the SNR response in the absence of the hard cutoff frequency at GHz, it was verified that this system can potentially achieve Gb/s at a pre-FEC BER of . Therefore, more work is yet to be done to realize the full potential of the SM MA VCSEL for ultra-high speed OWC system design.
Appendix A Performance Bounds of MMF Under Dispersion
The maximum bandwidth and reachable distance of MMF links under dispersion are derived, assuming Gaussian impulse responses for both CD and MD [17]. The RMS delay spread due to CD is given by [17]:
where is the CD coefficient in ps/nmkm, is the RMS spectral width of the laser source in nm, and is the fiber length in km. The dB cutoff frequency for the CD-induced pulse broadening is obtained as [17]:
The dB cutoff frequency corresponding to the MD-induced pulse broadening is [27]:
where the EMB of MMF cables is typically given in MHzkm. The RMS delay spread associated with MD is then given by:
The total RMS delay spread under the joint effect of CD and MD can be expressed as [17]:
which leads to the combined dB cutoff frequency of:
By substituting (9) and (10) into (13), the maximum bandwidth supported by MMF for a running distance of is derived as:
The maximum reachable distance is determined by the upper bound of such that based on (14), with representing the signal bandwidth, resulting in:
Acknowledgement
This work was financially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grant EP/Y037243/1 ‘Platform Driving The Ultimate Connectivity (TITAN)’.